A grocer paid $360 for some crates of fruit. Four crates spoiled and have to be thrown away, but he sold the rest at an increase in price of $2 a crate, gaining $88. How many crates did he buy?
Answer by mananth(16946)

assume he bought x crates
4 crates bad
so he sold (x-4) crates

cost of each crate = 360/x
he sold at ((360/x) +2 )
selling price - cost price = 88
((360/x) +2)*(x -4)-360 = 88
((360+2x)/x)*(x-4)-360 =88
multiply by x
(360+2x)(x-4)-360x=88x
360x-1440+2x^2-8x-360x=88x
2x^2-96x-1440=0
/2
x^2-48x-720=0
x^2-60x+12x-720=0
x(x-60)+12(x-60)=0
(x-60)(x+12)=0
x=60 OR -12
so he bought 60 crates